# Do LLMs have Consistent Values?

### v1

@@ Abstract
Values are a basic driving force underlying human behavior. Large Language Models (LLM) technology is constantly improving towards human-like dialogue. However, little research has been done to study the values exhibited in text generated by LLMs. Here we study this question by turning to the rich literature on value structure in psychology. We ask whether LLMs exhibit the same value structure that has been demonstrated in humans, including the ranking of values, and correlation between values. We show that the results of this analysis strongly depend on how the LLM is prompted, and that under a particular prompting strategy (referred to as “Value Anchoring”) the agreement with human data is quite compelling. Our results serve both to improve our understanding of values in LLMs, as well as introduce novel methods for assessing consistency in LLM responses.
@@ 1 Introduction
This question has only recently begun to be addressed. For example [1] show how probing LLMs with different names leads to variability which in some cases agrees with that of human populations. Here our focus is on understanding whether an LLM in a single conversation can exhibit a psychological characteristic profile that is similar to that of humans. This is a highly challenging question, since it requires analyzing a complete conversation and evaluating whether it conceivably could have been generated by a single individual.
@@
In order to stand on more quantitative ground to evaluating the quality of output relative to human research, we turn to the well established field of value psychology. Namely, we aim to quantify the values that LLM responses are aligned with, and whether these are in agreement with the value hierarchy and structure observed in humans. The question of values in LLMs has rarely been studied, and is of naturally broad interest. As an example of recent work, [2] prompt an LLM with a description of a profile of an individual characterized by a value and check whether generated text is consistent with this description. Our focus is very different, and asks whether an LLM response is in agreement with what we expect human responses to look like given research in the field.
@@
Values are basic motivations that play a foundational role in psychology, influencing perceptions and behaviors across various domains [3, 4], and representing fundamental aspects of human personality [5]. They have been extensively studied across diverse populations, and research has consistently demonstrated their enduring influence over behavior across time and contexts [4].
@@
One prominent framework for studying values, the Theory of Basic Human Values [6], outlines 19 core values, categorized by motivational goals [7]. These values can be simplified into a two-dimensional structure: conservation vs. openness to change, and self-enhancement vs. self-transcendence. The theory describes interrelations among values, suggesting that motivations driving some values are compatible with those driving other values, yet conflict with those underlying yet others. For instance, pursuing independence and creativity (self-direction) aligns with seeking change and variability (stimulation), but conflicts with an emphasis on the status quo (conformity). See Figure 1 for the theorized circle. Despite individual differences, there is a universal hierarchy, where for example caring for close others ranks high, while values related to dominance hold less importance across societies [8]. There is also ample empirical evidence that compatible values do tend to co-occur in humans [9, 10, 11], and are thus a “marker” of a human-like value system.
@@
Figure 1: Circular motivational continuum of 19 values in the refined value theory. Source: [12]. A value aligns with values that are adjacent on the circle and conflicts with those opposite to it. For example, self direction aligns with stimulation, and both conflict with conformity.
Our key question is therefore whether LLM responses demonstrate the same statistical behavior observed in humans: value-ranking and value-correlations. Note that this question is only meaningful when observing a single session of an LLM, where it can exhibit multiple values, because different sessions may be expressing different LLM “personas”. To study this question, we present LLMs with a value questionnaire (the Portrait Value Questionnaire—Revised – PVQ-RR– from [13], a well-established measure of values), and prompt them to answer all the questions in a single session (i.e., in the same context window). We then analyze the provided answers, putting specific emphasis on the correlation between answers in the same session.
We analyze two of the most recent large language models (LLMs): GPT 4 and Gemini Pro [^1] Our analysis also included GPT-3.5 and Palm2, which produced qualitatively similar results.. Our examination reveals that standard prompting of LLMs does not result in a population of human-like personas. We go on to explore prompting the LLMs with other prompts that provide additional information about the LLM persona. In particular we consider names [1] and persona descriptions. In addition, we consider a novel prompt which we refer to as a “Value Anchor”, which instructs the language model to answer as a person emphasizing a given value. We find that with these prompts, and in particular with the Value Anchor prompt, the overall statistics of the LLM responses (across sessions and within a session) closely mirror those of human subjects. Perhaps most surprising is our finding that the correlation between values agrees with the well known Schwartz circular model for correlations between values.
Our main findings are thus that LLMs cannot be considered as having a consistent, human like, value system that can be studied as representative of the values of the LLM. However, with appropriate prompting, LLMs can exhibit various coherent value personas similar to that of a population of humans. In other words, they answer a set of questions in a way that is consistent with a response from multiple human subjects. We provide information for best prompts and settings for this aim. In addition, we provide two datasets comprising 300 personas each, generated by GPT 4 and Gemini Pro. In conclusion, our results demonstrate the utility of using psychological theory for evaluating consistency of personas generated by LLMs.
@@ 2 Related Work
Values in LLMs: Our work is based on the Schwartz theory of Personal Values, a highly accepted theory within personality psychology [3]. Values are abstract goals, defining the end states individuals aspire for (e.g., safety, independence), used to direct judgements and behaviors [6, 7]. Individuals typically prioritize their values, so that values stemming from compatible motivations are similarly important, while values stemming from conflicting motivations are prioritized differently. These associations were replicated across hundreds of samples, across the world [14, 15], and make value theory especially useful to identify the coherence of the value profiles created by LLMs. Several studies assumed LLMs can be characterized as operating on the basis of a single set of values, taking an “LLMs as individuals” approach. Fischer and colleagues [2] tested whether ChatGPT could comprehend human values by providing it with value-related prompts and analyzing whether its responses matched the intended value category. A second study [16], compared ChatGPT’s values to those observed in the World Value Survey, while another [17] investigated how temperature influences GPT-3’s responses to the Human Value Scale. Scherrer and colleagues [18] studied responses of LLMs prompts evaluating moral positions, especially in ambiguous settings. A recent study [19], tested the value-like constructs embedded in LLMs and revealed both similarities and differences between LLMs and humans’ values. Kovak et al. [20] challenged those studies by establishing that context starkly influences values expressed by ChatGPT. They found significant variability in ChatGPT’s value expression in response to contextual changes, threatening the notion of stable characteristics of LLMs. Building upon the insights in [20], our study posits that upon providing controlled variability of context, LLMs can elicit a population of multiple personas. In this regard, we aim to further explore the accuracy of LLMs’ mimicking abilities within a controlled experimental framework.
@@
Prompting LLMs: There is extensive research on the prompt design for mimicking individual characteristics in LLM responses [21]. Approaches use specific scenarios [22], questionnaire items [23], simulation of social identities or areas of expertise [24], utilization of titles and surnames representing genders and ethnicities [1], and other demographic information [25]. Additionally, researchers explored the use of designated personas [26], and employed RLHF [27] to guide LLMs to reflect distinct personality traits. Despite this extensive body of work, to our knowledge, no study has directly compared the various prompting techniques to determine which approach yields responses that simulate within-session psychological characteristics of an individual best.
@@
Temperature in LLMs: Adjusting the temperature stands as a common practice for introducing variability in LLM responses [17]. However, consensus is lacking on the optimal temperature setting in simulating psychological characteristics. Some researchers advocate for higher temperatures to boost creativity [24], yet this can also introduce more noise into the data [28]. Conversely, setting the temperature to zero minimizes variability, enhancing replicability [27], albeit posing challenges for variance-dependent analysis [29]. Our framework enables us to explore how temperature adjustments impact the ability of LLMs to simulate human characteristics across multiple datasets.
@@
Evaluating the Quality of Persona Generation in LLMs: The ability of LLMs to mimic and portray human characteristics is a focus of intense research [30, 31]. LLMs can express psychological traits and attributes similar to human individuals [27, 32], and even simulate diverse populations [33, 24]. However, we are only beginning to understand the coherence of these LLM-generated characteristics in mirroring human psychological profiles [1, 20], and how to reliably produce such responses. We are specifically challenged to evaluate the coherence of the resulting psychological profiles. The literature suggested a number of approaches, including an open-ended interview with LLM-generated personas in order to assess the consistency between their intended characters and the responses [34]. In addition, one may apply an additional “judge” LLM in order to check an LLM persona [35]. Finally, [23] assessed coherence with a description used to prompt the LLM. Our study extends upon this line of research by applying well established characteristics of human psychology to investigate the quality of LLM generated personas.
@@ 3 Method
The Value Questionnaire: Our key goal was to assess responses of LLMs to questionnaires used to measure values in human subjects. Specifically, we considered the commonly used 57-item Portrait Value Questionnaire—Revised (PVQ-RR; [13]), developed to measure the 19 values in the Schwartz’s theory. The questionnaire describes fictional individuals and what is important to them. For example: “It is important to him/her to take care of people he/she is close to” (an item measuring benevolence-care values). For each such item, the subject is requested to indicate on a 6-point scale to what degree the persona they form is similar to the person described. Answers are categorical and range from a value of 1 (indicating “not like me at all”) to 6 (indicating “very much like me”). See Appendix for instructions and more example items from the questionnaire.
Models Used: We employed two prominent LLMs, specifically OpenAI’s GPT-4 and Google’s Gemini Pro. Each model was prompted with the five prompts (see Section 3.1), 300 times overall. Half of the runs applied the male-version of the questionnaire, and half the female version. The entire process was conducted twice, once with the temperature parameter set to $0.0$ and once with it set to $0.7$, resulting in the generation of 20 datasets for analysis.
@@ 3.1 Prompts
As mentioned above, we would like to measure the response of LLMs to PVQ-RR. However, as with many other LLM applications, the way the model is prompted has a significant effect on output. The instructions of the PVQ questionnaire were adjusted to suit LLMs. In cases of ambiguous gender identification in the prompt, we randomly allocated the female or male version to maintain consistency of administration. LLMs were then prompted to assess their likeness to the 57 descriptions incorporated in the PVQ-RR. They were instructed not to provide additional explanations to maintain focus on the task at hand. Following each prompt, the LLM was provided with all 57 items of the questionnaire in one administration. The study utilized a basic prompt, as well as four different prompts below that vary instructions to create multiple personas.
@@
Value Anchor prompt: This prompt adds an anchor of value importance using identification with an item used in an additional value questionnaire, akin to the approach outlined in the study by [23]. Participants are instructed as follows: “For each of the following descriptions, please answer how much the person described is like you from 1 (Not like me at all) to 6 (Very much like me), without elaborating on your reasoning. Answer as a person that is [value]”. Here “[value]” is taken from the Best-Worst Refined Values scale [36]. As a result, the prompts refer conceptually to the same values that are measured using the PVQ-RR, yet do not refer directly to the value items to be answered in response to the prompt. Examples of these anchor items include “protecting the natural environment from destruction or pollution” (universalism-nature) or “obeying all rules and laws” (conformity-rules). Please refer to E in the appendix for the complete list of anchor items.
@@
Demographic prompt: Drawing from the methodology of [25], this prompt extends the original prompt by incorporating additional demographic details. LLMs are asked to provide ratings based on the following prompt: “For each of the following descriptions, please rate how much the person described is like you, using a scale from 1 (Not like me at all) to 6 (Very much like me), without elaborating on your reasoning. Answer as a [age]-year-old who identifies as [gender], working in the field of [occupation], and enjoys [hobby].” The age, gender, occupation and hobby were randomly allocated for each prompt from a predefined list or range. The age range specified was between 18 and 75, with gender options including male, female, non-binary, and other, adapted from the National Academies of Sciences, Engineering, and Medicine [37]. Occupations were sourced from the World Values Survey (WVS-7; [38]), while hobbies were chosen from established lists supplied by The Activity Card Sort (ACS-UK; [39]). The lists of occupations and hobbies are available upon request.
@@
Generated Persona prompt: In line with the methodology of [40], we directed the models to craft personas. Our instruction was formulated as: “Create a persona (2-3 sentences long):”, with the temperature set at $0.7$ to stimulate the models’ creativity. An example of a persona generated by Gemini Pro is as follows: “Emily is a 25-year-old marketing manager who is passionate about her career and loves spending time with her friends and family. She is always looking for new ways to improve her skills and knowledge, and she is always up for a challenge.” Using these generated personas, we subsequently prompted the model as follows: “For each of the following descriptions, please rate how much the person described is like you, using a scale from 1 (Not like me at all) to 6 (Very much like me), without elaborating on your reasoning. Answer as: [persona].
@@
Names prompt: In line with a study by [1], the prompts comprised titles (i.e., Mr., Ms., and Mx.) followed by surnames representing five distinct ethnic groups. From the 500 names cataloged in the previous study, we randomly generated 300 unique combinations of titles and names, including 60 from each ethnic group. The prompt was structured as follows: “For each of the following descriptions, please rate how much the person described is like you, using a scale from 1 (Not like me at all) to 6 (Very much like me), without elaborating on your reasoning. Answer as [title + name]”. The complete list of titles and names are available upon request.

### v3

@@ Abstract
Large Language Models (LLM) technology is constantly improving towards human-like dialogue. Values are a basic driving force underlying human behavior, but little research has been done to study the values exhibited in text generated by LLMs. Here we study this question by turning to the rich literature on value structure in psychology. We ask whether LLMs exhibit the same value structure that has been demonstrated in humans, including the ranking of values, and correlation between values. We show that the results of this analysis depend on how the LLM is prompted, and that under a particular prompting strategy (referred to as “Value Anchoring”) the agreement with human data is quite compelling. Our results serve both to improve our understanding of values in LLMs, as well as introduce novel methods for assessing consistency in LLM responses.
@@ 1 Introduction
This question has only recently begun to be addressed. For example Aher et al. (2023) show how probing LLMs with different names leads to variability which in some cases agrees with that of human populations. Here our focus is on understanding whether an LLM in a single conversation can exhibit a psychological characteristic profile that is similar to that of humans. This is a highly challenging question, since it requires analyzing a complete conversation and evaluating whether it conceivably could have been generated by a single individual.
@@
In order to stand on more quantitative ground to evaluating the quality of output relative to human research, we turn to the well established field of value psychology. Namely, we aim to quantify the values that LLM responses are aligned with, and whether these are in agreement with the value hierarchy and structure observed in humans. The question of values in LLMs has rarely been studied, and is of naturally broad interest. As an example of recent work, Fischer et al. (2023) prompt an LLM with a description of a profile of an individual characterized by a value and check whether generated text is consistent with this description. Our focus is very different, and asks whether an LLM response is in agreement with what we expect human responses to look like given research in the field.
@@
Values are basic motivations that play a foundational role in psychology, influencing perceptions and behaviors across various domains (Sagiv and Schwartz, 2022, Sagiv et al., 2017), and representing fundamental aspects of human personality (Roberts and Yoon, 2022). Research has consistently demonstrated their enduring influence over behavior across time and contexts (Sagiv et al., 2017).
@@
One prominent framework for studying values, the Theory of Basic Human Values (Schwartz, 1992), outlines 19 core values, categorized by motivational goals (Schwartz, 2012). These values can be simplified into a two-dimensional structure: conservation vs. openness to change, and self-enhancement vs. self-transcendence. The theory describes interrelations among values, suggesting that motivations driving some values are compatible with those driving other values, yet conflict with those underlying yet others. For instance, pursuing independence and creativity (self-direction) aligns with seeking change and variability (stimulation), but conflicts with an emphasis on the status quo (conformity). See Figure 1 for the theorized circle. One of the key aspects of the theory is its cross-cultural coverage: it was developed to apply across populations, and tested in nearly 100 countries across all continents of the world, identifying points of commonality and differences among these populations (Sagiv and Schwartz, 2022). The use of the theory allows a stable and extremely general baseline of human values to compare LLMs. Hundreds of samples demonstrate individual differences in value importance. Importantly, they also demonstrate a universal hierarchy, where people are more likely to stress some values over others, notwithstanding the existing variability. For example caring for close others ranks high, while values related to dominance hold less importance across societies (Schwartz and Bardi, 2001). There is also ample empirical evidence that compatible values tend to be correlated in humans (Skimina et al., 2021a, Daniel et al., 2023, Schwartz and Cieciuch, 2022), and are thus a “marker” of a human-like value system. To summarize the above, human data pertains to both first-order statistics of values (i.e., which values rank high or low across the population), and second-order statistics (i.e., how are different values within an individual correlated).
@@
Figure 1: Circular motivational continuum of 19 values in the refined value theory. Source: Schwartz et al. (2012). A value aligns with values that are adjacent on the circle and conflicts with those opposite to it. For example, self direction aligns with stimulation, and both conflict with conformity.
Our key question is therefore whether LLM responses demonstrate the same statistical behavior observed in humans with respect to both value-ranking and value-correlations. Note that the question of value-correlations is of particular interest, because it allows bench-marking the extent to which responses of an LLM demonstrate a coherent “persona”. For example, while it is possible for a person to give a high score to Power Dominance, that person is unlikely to give a high score to Benevolence, since these are contradicting values.
To study this question quantitatively, we present LLMs with a value questionnaire (the Portrait Value Questionnaire—Revised – PVQ-RR– from Schwartz (2017), a well-established measure of values), and prompt them to answer all the questions in a single session (i.e., in the same context window). We then analyze the provided answers, putting specific emphasis on the correlation between answers in the same session.
We analyze two recent LLMs: GPT 4 and Gemini Pro, as well as four open models: Llama 3.1 8B, Llama 3.1 70B, Gemma 2 9B, and Gemma 2 27B. [^1] Our analysis also included GPT-3.5 and Palm2, which produced qualitatively similar results. Our results show that standard prompting of LLMs does not result in a population of human-like personas. We go on to explore prompting the LLMs with other prompts that provide additional information about the LLM persona. In particular we consider names (Aher et al., 2023) and persona descriptions. In addition, we consider a novel prompt which we refer to as a “Value Anchor”, which instructs the language model to answer as a person emphasizing a given value. We find that with these prompts, and in particular with the Value Anchor prompt, the overall first and second order statistics of the LLM responses closely mirror those of human subjects. Perhaps most surprising is our finding that the correlation between values agrees with the well known Schwartz circular model for correlations between values. We furthermore provide an explanation for how this correlation comes about. We provide information for best prompts and settings for this aim. In addition, we include six datasets comprising 300 personas each, generated by the models. In conclusion, our results demonstrate the utility of using psychological theory for evaluating consistency of personas generated by LLMs.
@@ 2 Related Work
Values in LLMs: Our work is based on the Schwartz theory of Personal Values, a highly accepted theory within personality psychology (Sagiv and Schwartz, 2022). Values are abstract goals, defining the end states individuals aspire for (e.g., safety, independence), used to direct judgements and behaviors (Schwartz, 1992; 2012). Individuals typically prioritize their values, so that values stemming from compatible motivations are similarly important, while values stemming from conflicting motivations are prioritized differently. These associations were replicated across hundreds of samples, across the world (Pakizeh et al., 2007, Skimina et al., 2021b), and make value theory especially useful to identify the coherence of the value profiles created by LLMs. Several studies assumed LLMs can be characterized as operating on the basis of a single set of values, taking an “LLMs as individuals” approach. Fischer et al. (2023) tested whether ChatGPT could comprehend human values by providing it with value-related prompts and analyzing whether its responses matched the intended value category. A second study Lindahl and Saeid (2023), compared ChatGPT’s values to those observed in the World Value Survey, while another Miotto et al. (2022) investigated how temperature influences GPT-3’s responses to the Human Value Scale. Scherrer and colleagues Scherrer et al. (2023) studied responses of LLMs prompts evaluating moral positions, especially in ambiguous settings. A recent study by Hadar-Shoval et al. (2024), tested the value-like constructs embedded in LLMs and revealed both similarities and differences between LLMs and humans’ values. Kovač et al. (2023) challenged those studies by establishing that context starkly influences values expressed by ChatGPT. They found significant variability in ChatGPT’s value expression in response to contextual changes, threatening the notion of stable characteristics of LLMs. Building upon the insights in Kovač et al. (2023), our study posits that upon providing controlled variability of context, LLMs can elicit a population of multiple personas. In this regard, we aim to further explore the accuracy of LLMs’ mimicking abilities within a controlled experimental framework.
@@
Prompting LLMs: There is extensive research on prompt design for mimicking individual characteristics in LLMs (Liu et al., 2023). Approaches use specific scenarios (Hadar-Shoval et al., 2023), questionnaire items (Jiang et al., 2023), simulation of social identities or areas of expertise (Salewski et al., 2024), utilization of titles and surnames representing genders and ethnicities (Aher et al., 2023), and other demographic information (Argyle et al., 2023). Additionally, researchers explored the use of designated personas (Safdari et al., 2023), and employed RLHF (Li et al., 2023) to guide LLMs to reflect distinct personality traits. Despite this extensive body of work, to our knowledge, no study has directly compared the various prompting techniques to determine which approach yields responses that simulate within-session psychological characteristics of an individual best.
@@
Temperature in LLMs: Adjusting the temperature stands as a common practice for introducing variability in LLM responses (Miotto et al., 2022). However, consensus is lacking on the optimal temperature setting in simulating psychological characteristics. Existing research includes use of mostly two temprature settings: 0.7 and 0. Some researchers advocate for higher temperatures to boost creativity (Salewski et al., 2024), yet this can also introduce more noise into the data (Gunel et al., 2020). Conversely, setting the temperature to zero minimizes variability, enhancing replicability (Li et al., 2023), albeit posing challenges for variance-dependent analysis (Hagendorff et al., 2023). Our framework enables us to explore how temperature adjustments impact the ability of LLMs to simulate human characteristics across multiple datasets.
@@
Evaluating the Quality of Persona Generation in LLMs: The ability of LLMs to mimic and portray human characteristics is a focus of intense research (Binz and Schulz, 2023, Ouyang et al., 2022). LLMs can express psychological traits and attributes similar to human individuals (Li et al., 2023, Stevenson et al., 2022), and even simulate diverse populations (Deshpande et al., 2023, Salewski et al., 2024). However, we are only beginning to understand the coherence of these LLM-generated characteristics in mirroring human psychological profiles (Aher et al., 2023, Kovač et al., 2023), and how to reliably produce such responses. We are specifically challenged to evaluate the coherence of the resulting psychological profiles. The literature suggested a number of approaches, including an open-ended interview with LLM-generated personas in order to assess the consistency between their intended characters and the responses (Wang et al., 2024). In addition, one may apply an additional “judge” LLM in order to check an LLM persona (Gupta et al., 2024). Finally, Jiang et al. (2023) assessed coherence with a description used to prompt the LLM. Our study extends upon this line of research by applying well established characteristics of human psychology to investigate the quality of LLM generated personas.
@@ 3 Method
The Value Questionnaire: Our key goal was to assess responses of LLMs to questionnaires used to measure values in human subjects. Specifically, we considered the commonly used 57-item Portrait Value Questionnaire—Revised (PVQ-RR; (Schwartz, 2017)), developed to measure the 19 values in the Schwartz’s theory. The questionnaire describes fictional individuals and what is important to them. For example: “It is important to him/her to take care of people he/she is close to” (an item measuring benevolence-care values). For each such item, the subject is requested to indicate on a 6-point scale to what degree the persona they form is similar to the person described. Answers are categorical and range from a value of 1 (indicating “not like me at all”) to 6 (indicating “very much like me”). See Appendix for instructions and more example items from the questionnaire.
Models Used: We employed six prominent LLMs, specifically OpenAI’s GPT-4, Google’s Gemini Pro, Llama 3.1 8B, Llama 3.1 70B, Gemma 2 9B, and Gemma 2 27B. Each model was prompted with the five prompts (see Section 3.1), 300 times overall. Half of the runs applied the male-version of the questionnaire, and half the female version. The entire process was conducted twice, once with the temperature parameter set to $0.0$ and once with it set to $0.7$, resulting in the generation of 20 datasets for analysis.
@@ 3.1 Prompts
As mentioned above, we would like to measure the response of LLMs to PVQ-RR. However, as with many other LLM applications, the way the model is prompted has a significant effect on output. The instructions of the PVQ questionnaire were similar to those in prior research, but with added text instructing the LLM not to elaborate. This resulted in the LLM producing only the value scores, thereby simplifying processing and analysis. LLMs were prompted to assess their likeness to the 57 descriptions incorporated in the PVQ-RR. Following each prompt, the LLM was provided with all 57 items of the questionnaire in one administration. [^2] We also conducted a serial prompting analysis for the Llama models. The results comparing batch administration to serial administration are detailed in the Appendix Figure 6 and Table 5 The study utilized a basic prompt, as well as four different prompts below that vary instructions to create multiple personas.
@@
Value Anchor prompt: This prompt adds an anchor of value importance using identification with an item used in an additional value questionnaire, akin to the approach outlined in the study by Jiang et al. (2023). Participants are instructed as follows: “For each of the following descriptions, please answer how much the person described is like you from 1 (Not like me at all) to 6 (Very much like me), without elaborating on your reasoning. Answer as a person that is [value]”. Here “[value]” is taken from the Best-Worst Refined Values scale (Lee et al., 2019). As a result, the prompts refer conceptually to the same values that are measured using the PVQ-RR, yet do not refer directly to the value items to be answered in response to the prompt. Examples of these anchor items include “protecting the natural environment from destruction or pollution” (universalism-nature) or “obeying all rules and laws” (conformity-rules). Please refer to E in the appendix for the complete list of anchor items.
@@
Demographic prompt: Drawing from the methodology of Argyle et al. (2023), this prompt extends the original prompt by incorporating additional demographic details. LLMs are asked to provide ratings based on the following prompt: “For each of the following descriptions, please rate how much the person described is like you, using a scale from 1 (Not like me at all) to 6 (Very much like me), without elaborating on your reasoning. Answer as a [age]-year-old who identifies as [gender], working in the field of [occupation], and enjoys [hobby].” The age, gender, occupation and hobby were randomly allocated for each prompt from a predefined list or range. The age range specified was between 18 and 75, with gender options including male, female, non-binary, and other, adapted from the National Academies of Sciences, Engineering, and Medicine (National Academies of Sciences, Engineering, and Medicine, 2022). Occupations were sourced from the World Values Survey (WVS-7; (Haerpfer et al., 2022)), while hobbies were chosen from established lists supplied by The Activity Card Sort (ACS-UK; (Laver-Fawcett et al., 2016)). The lists of occupations and hobbies are available upon request.
@@
Generated Persona prompt: In line with the methodology of Cheng et al. (2023), we directed the models to craft personas. Our instruction was formulated as: “Create a persona (2-3 sentences long):”, with the temperature set at $0.7$ to stimulate the models’ creativity. An example of a persona generated by Gemini Pro is as follows: “Emily is a 25-year-old marketing manager who is passionate about her career and loves spending time with her friends and family. She is always looking for new ways to improve her skills and knowledge, and she is always up for a challenge.” Using these generated personas, we subsequently prompted the model as follows: “For each of the following descriptions, please rate how much the person described is like you, using a scale from 1 (Not like me at all) to 6 (Very much like me), without elaborating on your reasoning. Answer as: [persona].
@@
Names prompt: In line with a study by Aher et al. (2023), the prompts comprised titles (i.e., Mr., Ms., and Mx.) followed by surnames representing five distinct ethnic groups. From the 500 names cataloged in the previous study, we randomly generated 300 unique combinations of titles and names, including 60 from each ethnic group. The prompt was structured as follows: “For each of the following descriptions, please rate how much the person described is like you, using a scale from 1 (Not like me at all) to 6 (Very much like me), without elaborating on your reasoning. Answer as [title + name]”. The complete list of titles and names are available upon request.